### v1

# Generative Emergent Communication: Large Language Model is a Collective World Model

Abstract
This study proposes a unifying theoretical framework called generative emergent communication (generative EmCom) that bridges emergent communication, world models, and large language models (LLMs) through the lens of collective predictive coding (CPC). The proposed framework formalizes the emergence of language and symbol systems through decentralized Bayesian inference across multiple agents, extending beyond conventional discriminative model-based approaches to emergent communication. This study makes the following two key contributions: First, we propose generative EmCom as a novel framework for understanding emergent communication, demonstrating how communication emergence in multi-agent reinforcement learning (MARL) can be derived from control as inference while clarifying its relationship to conventional discriminative approaches. Second, we propose a mathematical formulation showing the interpretation of LLMs as collective world models that integrate multiple agents’ experiences through CPC. The framework provides a unified theoretical foundation for understanding how shared symbol systems emerge through collective predictive coding processes, bridging individual cognitive development and societal language evolution. Through mathematical formulations and discussion on prior works, we demonstrate how this framework explains fundamental aspects of language emergence and offers practical insights for understanding LLMs and developing sophisticated AI systems for improving human-AI interaction and multi-agent systems.
Index Terms:
emergent communication, large language model, world model, generative model, symbol emergence.
I Introduction
Langauge evolves and changes over time as a result of decentralized human communications [1, 2, 3, 4]. Sentences are generated to describe a wide range of phenomena, including external events, emotions, and intentions. In particular, the system of language is not static but dynamic [5, 6, 7, 8]. As Peirce, the founder of semiotics, suggested, symbols, including language, can be characterized by a triadic relationship of sign, object, and interpretant [9, 10, 11]. Here, sign corresponds to words, sentences and other signals. In particular, the correspondence between sign and object, which is signified by a sign (i.e., signifier), is determined by an interpretant. In other words, the meaning of a sign, that is, language, depends on culture and context, and so on.
The series of studies on emergent communication (EmCom), which is also referred to as emergent language (EmLang), and symbol emergence has attempted to explain the emergence of language and sharing of meanings of language among agents [2, 12, 13, 14, 15, 16]. However, computational models and general theories that provide a comprehensive and integrative understanding of symbol and language emergence [17, 8] and address the interdependency between the two aspects are lacking: first, world modeling by agents, which depends on their embodiment and environmental adaptation [18, 19]; and second, language emergence, which involves the evolution of language structure reflecting structural knowledge of the world that becomes embedded in language through distributional semantics[20, 21]. Although we are entering the era of large language models (LLMs) and generative artificial intelligence (AI), a need for a theoretical framework that can explain the dynamic and semantic aspects of language emergence in embodied cognitive developmental systems still remains [22, 8, 23, 16].
Recently, numerous studies have discussed how LLMs possess a model of the world. LLMs learn the distribution of word or linguistic token sequences and become intelligent to the extent that they can solve a wide range of tasks, including question answering, machine translation, and conversations [24, 25]. These capabilities are widely believed to be based on distributional semantics [20, 22]. However, the emergence of a latent structure that enables distributional semantics in our human language is not yet uncovered, though the mechanism through which neural networks find the hidden structure of language is gradually being revealed. Even before the invention of LLMs, word2vec, skip-gram, or continuous bag of words language models could internally form relative knowledge of concepts, for instance, "$\rm{London}"-"\rm{UK}"+"\rm{France}"\simeq"\rm{Paris}"$ [26, 27]. Two language models independently modeling different languages can perform unsupervised machine translation through structural alignment to a certain extent [28]. Recently, it has become clearer that LLMs have much knowledge of the world, including color similarity, physical properties of materials, and spatial knowledge [29, 30, 31, 32, 33]. This implies that the capability of LLMs is based on the mysterious nature of language. We believe that uncovering this nature of human language is the crucial mission of studies on symbol/language emergence and EmCom/EmLang.
In related fields, for instance, evolutionary robotics, cognitive and developmental robotics, artificial life, computational linguistics, and machine learning (ML), a wide range of studies on language evolution, EmCom, and symbol emergence have been conducted [34, 14, 2, 23, 17]. Most of these studies relied on specific types of language games, for instance, referential, signaling, and naming games. Since the late 2010s, empowered by the representation learning and language modeling capability of deep neural networks (DNNs), studies on EmCom have been boosted again [35, 36, 37]. However, most studies only treated the formation of communication protocols in specific game settings.
These approaches somehow failed to construct a general framework capturing symbol emergence from the viewpoint perspective of general principles of environmental adaptation, such as the free-energy principle (FEP), predictive coding (PC), and world modeling. Recently, world models have garnered significant attention as representation-learning models that incorporate action outputs and temporal dynamics of agent– environment interactions [38, 18, 19]. This aligns with broader theoretical frameworks such as PC and the FEP. PC posits that the brain constantly predicts sensory information and updates its internal models to enhance predictability  [39], whereas FEP provides a more generalized framework explaining the self-organization of biological systems through minimization of free energy [40], which is associated with the idea of the Bayesian brain proposed by [41]. Notably, FEP extends beyond individual cognition to explain the self-organization of cognitive and biological systems in detail [42, 43, 44], making it a promising foundation for understanding symbol emergence at both individual and collective levels.
To address these challenges, Taniguchi proposed the collective predictive coding (CPC) hypothesis [16]. The idea is based on a generative view of cognitive systems. The CPC assumes that not only individual agents but also groups of agents committing to symbolic communications can be modeled as generative models. Regarding individual agents, the FEP and active inference provide a generative view of a cognitive system. In other words, the CPC hypothesis extends the idea of the FEP and PC to a societal level [45]. The CPC hypothesis argues that symbol/language emergence can be modeled as decentralized Bayesian inference of a shared latent representation in a hierarchical manner.
This study aims to provide a theoretical view that bridges the idea of world models, that is, internal models for the sensorimotor observations and dynamics, and the emergence of language. To this end, we extend the idea of the CPC hypothesis and formalize the framework of generative emergent communication (generative EmCom) by introducing a generative view to the existing framework of EmCom. This allows us to develop a more integrative general theory of multi-agent systems performing EmCom and symbol emergence systems.
The main contributions of this paper are as follows:
  1. 1.
    Propose a new framework of EmCom known as generative EmCom, demonstrating that the emergence of communication in multi-agent reinforcement learning (MARL) can be derived from the perspective of control as inference (CaI) within the framework of generative EmCom, and clarify the relationship between conventional discriminative EmCom and the new generative EmCom.
  2. 2.
    Present the mathematical framework showing how LLMs become a collective world model in a specific sense through collective PC.
The remainder of this paper is organized as follows: Section 2 introduces the framework of generative EmCom, formalizing the collective PC hypothesis and its mathematical foundations. Section 3 demonstrates the formulation of language games as decentralized Bayesian inference, with a focus on the Metropolis-Hastings naming game (MHNG). Section 4 explores the application of generative EmCom to MARL, showing how it facilitates cooperative behavior. Section 5 discusses the relationship between the proposed generative framework and conventional approaches to EmCom. Section 6 examines LLMs through the lens of collective world models, providing new insights into their capabilities and limitations. Finally, Section 7 concludes the paper and discusses future directions for research in this area.
II Generative EmCom
Figure 1: (Left) PGM representation of CPC in symbol emergence, that is, emergent communication or language. The top-down generative process corresponds to language understanding and prediction of observations, which is downward causation in the symbol emergence system. The bottom-up inference process corresponds to perception, representation learning, and communication, that is, language game, which is upward causation in the symbol emergence system. (Right) Overview of the CPC in a symbol emergence system illustrating the bidirectional process of language understanding and generation, mediated by inference through language games and representation learning.
II-A Collective predictive coding
The CPC hypothesis was proposed to explain the emergence of symbol systems, particularly language [16, 46]. The CPC extends the idea of PC and FEP from the individual cognition level to a societal level. Taniguchi et al. were the first to formulate the CPC from the perspective of FEP and active inference [45]. This suggests that symbol emergence follows a process of minimizing free energy across the entire multi-agent system.
Although PC theory suggests that individual brains constantly predict sensory information and update their internal representations, including world models. CPC suggests that a group of agents, for instance, a human society, predict sensory information of all of the agents and update its external representations, that is, symbol systems.
A question is raised. How can we update the external representations, e.g., language, while our brains are disconnected physically? The CPC hypothesis suggests that a type of language game performs a decentralized Bayesian inference among the group (e.g., [47, 48]). In this framework, language games (such as naming games) can be interpreted as implementing decentralized Bayesian inference of shared representations. A representative example is MHNG explained in Section II-C. The CPC hypothesis argues that symbol systems emerge as a result of decentralized Bayesian inference performed collaboratively by multiple agents.
Although the encoding of sensory information through internal representations is ensured by the plasticity of neural systems, the plasticity of external representations is guaranteed by the flexibility of our symbol systems. The arbitrariness of symbol systems is a widely recognized characteristic of symbols in semiotics [11]. Peirce referred to the process by which subjects assign meaning to symbols according to culture and context as the semiosis. Although our brains are physically and electrically separated, they are informationally connected through communication using a flexible symbol system. Therefore, with appropriate communication and symbol system update algorithms, we can encode information into the symbol system as an external representation. In fact, the CPC hypothesis can consider that humans collectively perform this action in language emergence.
This implies that language collectively encodes information about the world as observed by numerous agents through their sensory-motor systems. The CPC hypothesis study [16] did not provide a clear and detailed explanation regarding this point while proposing a new perspective on why LLMs seem to possess knowledge about the real world. This is one of the main topics of this paper.
Essentially, CPC hypothesizes that human language is formed through a process of collective PC, where the symbol system emerges to maximize the predictability of multi-modal sensory-motor information obtained by members of a society, that is, minimize the collective free energy of a group of agents. This approach provides a unified framework for understanding symbol emergence, language evolution, and the nature of linguistic knowledge from the perspective of environmental adaptation and brain science.
II-B Generative EmCom
Numerous studies on EmCom/EmLang have been conducted. Major approaches are language games, MARL, and iterated learning models (ILMs) [12, 6, 49, 50, 51]. Recently, EmCom studies based on language games, particularly referential games, have become a dominant approach in the research community of EmCom. These studies utilize the representation learning capability of DNNs, give agents language encoders and decoders, and make them learn to generate and interpret a complex sequence of tokens to identify its meaning. In a typical setting of referential games, a speaker agent encodes a given object to a sequence of tokens using a neural network, for instance, long short-term memory, and the listener agent decodes it and identifies the object among some other items. This approach considers EmCom as an optimization of communication protocol. Generally, this is in line with Shannon’s communication model [52]. This approach focuses less on the synergy between the perceptual and communicative aspects of language. This approach can be regarded as a discriminative model-based approach to EmCom/EmLang.
Unlike traditional EmCom models, generative EmCom models consider language emergence as a generative model. In the simplest case, a probabilistic graphical model (PGM) of generative EmCom based on CPC hypothesis can be described mathematically as follows:
Generative model:
$p\left(\{x^{k}\}_{k},\{z^{k}\}_{k},m\right)=p(m)\prod_{k}p\left(x^{k}\mid z^{k}\right)p\left(z^{k}\mid m\right)$ (1)
Inference model:
$q\left(m,\{z^{k}\}_{k}\mid\{x^{k}\}_{k}\right)=q\left(m\mid\{z^{k}\}_{k}\right)\prod_{k}q\left(z^{k}\mid x^{k}\right)$ (2)
where $p(\cdot)$ and $q(\cdot)$ are assumed to be original probability density functions and approximate probability functions, respectively, following the conventional notation of variational inference. Variables and functions used in this paper are listed in Table I. In this framework, $q\left(z^{k}\mid x^{k}\right)$ is interpreted as the process of representation learning performed by agent $k$, whereas $q\left(m\mid\{z^{k}\}_{k}\right)$ is hypothesized to be derived through collaborative language games among the agents. The overall phenomenon of symbol emergence within the group can be considered as a distributed effort to approximate $q\left(m,\{z^{k}\}_{k}\mid\{x^{k}\}_{k}\right)$ without centralized control.
Mathematically, the CPC frames the process of symbol emergence in terms of social representation learning, analogous to the formation of internal representations by individual brains. As we introduced in Section 3, the representation learning and the inference of latent variables can be naturally extended to multi-agent reinforcement learning (RL) involving EmCom. Additionally, the formulation of CPC from the perspective of active inference can make us understand that this generative EmCom is not only for creating communication protocol but also for the environmental adaptation as a system of agents. Similarly, FEP and active inference are the general explanatory theory of the brain’s adaptation to the environment. In summary, the framework of generative EmCom based on the CPC hypothesis offers an explanation for language evolution to serve environmental adaptation at a collective level.
Generative EmCom offers an intriguing dual perspective, depending on whether one focuses on the collective or the individual. From a holistic perspective, the system appears to perform centralized representation learning of $w$ across all agents. Notably, $o^{k}$ can be easily extended to include multiple sensory modalities. From the perspective of the system overseer, or generative EmCom, multi-agent symbol emergence becomes analogous to multi-modal representation learning of external representations. Conversely, from the perspective of an individual agent, the system engages in multi-modal representation learning, influenced by and contributing to the inference of a shared latent variable $m$, which parameterizes the prior distribution common to all agents. The equivalence of these two interpretations is the cornerstone of the generative EmCom perspective. In essence, generative EmCom can be formulated as the representation learning of shared external representations, mediated through the representation learning of internal representations by each agent, based on multi-agent and multi-modal sensory-motor information.
TABLE I: Nomenclature and parameter details
$m_{t}$ Message (shared latent variable) communicated between agents at time $t$
$o^{k}_{t}$ Optimality variable for the $k$-th agent (1: optimal, 0: not optimal)
$z^{k}_{t}$ Internal representation (e.g., state in RL) of the $k$-th agent at time $t$
$a^{k}_{t}$ Action of the $k$-th agent at time step $t$
$x^{k}_{t}$ Observations (or sensory inputs) of the $k$-th agent at time $t$
$\theta^{k}$ Global parameters of the internal models of the $k$-th agent, for instance, neural networks
$\phi^{k}$ Parameters of language model of the $k$-th agent.
$r^{k}$ Reward function for the $k$-th agent
$p(\cdot)$ Original probability density function (i.e., a generative model)
$q(\cdot)$ Approximate probability function (i.e., an inference model)
II-C Language Game as Decentralized Bayesian Inference
The hypothetical argument that language game can perform the decentralized Bayesian inference has a computational basis though whether actual language communication can realize such decentralized Bayesian inference in our human society is an open question. The MHNG is an instance of this idea.
The MHNG comprises the following steps:
  1. 1.
    Perception: Speaker and listener agents (Sp and Li) observe the $d$-th object, obtain $x^{Sp}_{d}$, and $x^{Li}_{d}$ infers their internal representations $z_{d}^{Sp}$ and $z_{d}^{Li}$, respectively.
  2. 2.
    MH communication: Speaker mentions the name $m_{d}^{Sp}$ of the $d$-th object by sampling it from $P(m_{d}|z_{d}^{Sp},\phi^{Sp})$. The listener determines whether it accepts the naming with probability $\gamma={\rm min}\left(1,\frac{P(z_{d}^{Li}|{\phi}^{Li},m_{d}^{Sp})}{P(z_{d}^{Li}|{\phi}^{Li},m_{d}^{Li})}\right)$.
  3. 3.
    Learning: After MH communication was performed for every object, the listener updates its global parameters $\theta^{Li}$ and $\phi^{Li}$.
  4. 4.
    Turn-taking: The speaker and listener alternate their roles and go back to (1).
It has been demonstrated that the MHNG is equivalent to the Metropolis-Hastings algorithm for inferring latent variables in a probabilistic generative model. This model conditions the internal representations $z^{k}$ of multiple agents, acting as representation learning machines, on a common external representation $m$. Although the original study assumed two agents and a categorical message $m$, the core PGM of generative EmCom underlying this theory does not make these assumptions. Consequently, this fundamental idea can be extended in various ways.
As described, unlike referential games, MHNG assumes a joint attention performed by two agents. This assumption may seem strange from the game-theoretic approach to EmCom, such as referential games. However, developmental studies suggest that joint attention is fundamental to human infant language acquisition. Human infants develop joint attention capabilities before vocabulary explosion occurs, and joint attention serves as a crucial foundation for language acquisition [7]. Gergely et al. highlighted that when children incorporate parental instructions into their learning process, they presuppose that parents have the intention to teach them [53].
Moreover, interactions modeled by referential games—where agents identify objects and receive feedback—are rarely observed in natural language acquisition in cognitive development. Unlike game-theoretic approaches, MHNG, as an example of generative EmCom, assumes cooperative behavior in symbol emergence, which aligns with scholarly observations about human language acquisition [7]. Thus, although conventional (discriminative) EmCom provides rational models for both cooperative and competitive games, the assumptions of generative EmCom better align with empirical findings in human communication behavior.
Figure 2: (Top) Overview of the MHNG process. In the game, which two agents (A and B) engage in, the agents perceive a target object with joint attention and form internal representations. One agent (speaker) utters a sign and the other (listener) determines whether to accept it. Thereafter, they take turns. (Bottom) PGM representation, which is assumed in the MHNG. The MHNG is proved to be an inference process in the representation learning in a collective multi-agent system [47].
The original idea of MH-based EmCom was introduced by Hagiwara et al. [48] and later generalized and formalized by Taniguchi et al. [47], who clarified its connection to representation learning. The concept of MHNG has since been extended and validated in various ways. Furukawa et al. [54] demonstrated that similar symbol emergence can be observed even when the relationship between $m$ and $z$ is head-to-head instead of tail-to-tail. The extension to multimodal sensory information was achieved by Hagiwara et al. [55], who demonstrated that MHNG can lead to symbol emergence even when agents have different sensory modalities. Their work also showed that modality information possessed by other agents, but not by oneself, can facilitate object category formation. Similar multimodal extensions have been explored with variational autoencoder (VAE)-based representation learning by Hoang et al. [56], though their work suggests that the mechanism for integrating multimodal sensory information influences symbol emergence.
Although these MHNG studies involve two agents, a mathematical extension to N-agent conditions was developed by Inukai et al. [57], who introduced a recursive structure in communication while maintaining its characteristics as distributed Bayesian inference. They argued that random partner selection for MHNG can be considered as a one-sample and limited-length approximation of this approach.
Although several generative EmCom studies define the sign $m$ as a categorical variable, this is not a necessary assumption of the framework. A recent study by Hoang et al. [58] demonstrates that sharing compositional word sequences is possible within the MHNG framework, similar to numerous EmCom studies. From Peirce’s semiotic perspective, signs include both compositional discrete sequences and continuous signs such as voice pitch and facial expressions. A recent study by You et al. [59] demonstrated that such continuous signs can emerge within the generative EmCom framework, whereas Saito et al. [60] modeled the emergence of compositional signs from continuous time-series information as signals.
To verify the CPC hypothesis, examining whether human sign acceptance rates in joint attention naming scenarios match MHNG predictions was necessary. An experimental semiotic study by Okumura et al. [61] demonstrated that the sign acceptance rate in MHNG effectively predicts human behavior.
MHNG serves as a basic language game to represent symbol emergence by realizing distributed Bayesian inference and inferring latent variables of the generative model, corresponding to language. The critical point is that the inference of the posterior distribution is performed in a decentralized manner through language games. This suggests that being based on the MH method is not a necessary condition. The MHNG-based approach represents an initial step in modeling symbol emergence (or language emergence, EmCom) as decentralized Bayesian inference. Future studies should explore the possibility of constructing generative EmCom models by distributing various inference methods (e.g., [62]).
III EmCom for Multi-agent Cooperation
III-A Generative EmCom for Multi-agent Reinforcement Learning
Figure 3: Graphical model of generating cooperative actions for two agents.
Communication and language are often considered to emerge to facilitate multi-agent cooperation. In recent years, studies on MARL with communication channels have been progressing. Initial methods in multi-agent deep RL include DIAL [63] and CommNet [64]. These methods connect the networks of agents through messages, enabling the learning of necessary messages for cooperative behavior through backpropagation. Additionally, multi-agent deep deterministic policy gradient (MADDPG), an extension of the DDPG [65] for MARL, has been proposed [66]. In these methods, the agents use messages sent from all other agents to learn their policies. Methods involving weighting or attention mechanisms have been developed to limit communication to only necessary agents [67, 68, 69, 70]. To compute efficient messages, graph neural networks (GNNs) are used [71, 72, 73, 74, 75]. In these studies, multiple agents are connected through a network and messages are inferred by making them differentiable variables through backpropagation. In other words, error information computed from the internal states of others is directly transmitted to oneself, which is an unnatural modeling from the perspective of communication among independent individuals.
In contrast, generative EmCom allows us to formulate MARL with emergent communication in a Bayesian manner by incorporating the idea of CaI [76], which is a theory to formulate RL as a PGM [77, 78]. Figure 3 shows a graphical model of cooperative action generation between two agents, and the details of each stochastic variable are listed in Table I. The behavior of each agent is generated through a Markov decision process with a prior variable $m_{t}$. The state $z^{k}_{t}$ of an agent at time $t$ is determined according to state $z^{k}_{t-1}$, action $a^{k}_{t-1}$, and message $m_{t}$, which is the shared latent variable:
$z^{k}_{t}\sim p(z^{k}_{t}|m_{t},z^{k}_{t-1},a^{k}_{t-1}).$ (3)
where $k\in\{A,B\}$ denotes an index of agents. The agent can indirectly infer the state of others through the message $m_{t}$ in a probabilistic manner.
The optimality variable $o^{k}_{t}\in\{0,1\}$ represents the state optimality of both agents: $1$ indicates that the state and action pair are on the optimal trajectories, whereas 00 indicates it is not. Note that optimality is a type of probabilistic interpretation of reward functions as shown below. In this model, two types of optimality exist: one for each individual agent and the other for the group of agents. The probability $p(o^{k}_{t}=1|z^{k}_{t},a^{k}_{t})$ of this optimality variable is computed using reward function $r^{k}(z^{k}_{t},a^{k}_{t})$ as follows:
$p(o^{k}_{t}=1|z^{k}_{t},a^{k}_{t})\propto\exp(r(z^{k}_{t},a^{k}_{t})).$ (4)
For the reward of the group of agents $r^{AB}$, the following approximate reward models are learned by each agent:
$p(o^{AB}_{t}=1|z^{k}_{t},a^{k}_{t})\approx\propto\exp(r^{AB}(z^{A}_{t},a^{A}_{t},z^{B}_{t},a^{B}_{t})).$ (5)
[…]
Following the theory of CaI, the optimal state sequence for both agents can be calculated by inferring state $z_{t}$ and message $m_{t}$ under the condition that the value of the optimality variables is always 1, as if the two-agent system acts as a single agent:
$z^{A}_{t},m_{t}\sim p(z^{A}_{t},m_{t}|z^{B}_{t},o^{A}_{1:T}={\mathbf{1}},o^{AB}_{1:T}={\mathbf{1}}).$ (6)
However, this equation has two problems: it includes the internal state $z^{B}_{t}$ of others, which cannot be observed in practice, and deriving this probability distribution analytically is difficult. We solve these problems by alternately inferring the following two variables:
[…]
$m_{1:T}\sim p(m_{1:T}|z^{A}_{1:T},z^{B}_{1:T},o^{AB}_{1:T}={\mathbf{1}}){\rm:communication.}$
Equation (7) describes state planning, which can be computed based on the CaI framework [76]. Equation (8) describes the inference of the message and can be formulated using the MHNG proposed by Taniguchi et al. [55, 47], which allows both agents to infer messages through communication without observing each other’s internal states. Note that MHNG enables two agents to perform sampling in (8) without simultaneous observations of $z^{A}_{t},z^{B}_{t}$.
Thus, MHNG can be used not only for multimodal object categorization and naming but also for action coordination among multi-agents using PGM to formulate MARL, that is, modeling EmCom for multi-agent cooperation.
III-B Generative EmCom on World models
The generative EmCom for MARL can be extended to involve representation learning in the same way as we discussed in Section 2.3. This extension implies the integration of the idea of world models into generative EmCom.
The concept of a world model represents an internal model within an agent that captures the dynamics of environmental states, their responses to the actions of the agent, and their relationships with sensory inputs [79, 18, 19]. The concept of world model has its origins in the early days of AI and robotics studies [80]. Initial studies on ML investigated techniques for agents to autonomously construct and adapt their world models [81, 82]. Currently, the term generally refers to predictive frameworks [83], predominantly implemented using DNN architectures.
The concept of world models is closely related to the idea of PCs. PC is the idea that the brain constantly predicts sensory information and updates its internal models to minimize prediction errors. In the context of cognitive robotics and AI, world models provide the structure for representing and reasoning about the environment, whereas PC offers a mechanism for learning and updating these models based on sensory experiences. The free energy principle and active inference further unify these concepts, suggesting that both perception and action can be considered as processes of minimizing prediction errors or free energy. A theoretical connection exists between them[19] .
Generally, a theory of world models can be based on partially observable Markov decision process (POMDP). In this framework, the state $z_{t}$ at time $t$ is not directly observable by the agent. Instead, the agent receives an observation $x_{t}$, which is assumed to be generated from a latent state $z_{t}$. The agent’s actions $a_{t}$ influence the transition of states according to a probability distribution $p(z_{t+1}|z_{t},a_{t})$. The observation model is given by $p(x_{t}|z_{t})$. The goal of the agent is to learn these probability distributions and use them to make predictions and inferences about the environment. This can be formalized as:
State transition: $\quad z_{t+1}\sim p(z_{t+1}|z_{t},a_{t})$ (9)
Observation: $\quad x_{t}\sim p(x_{t}|z_{t})$ (10)
Inference: $\quad z_{t}\sim q(z_{t}|x_{1:t},a_{1:t-1})$ (11)
where $q(z_{t}|x_{1:t},a_{1:t-1})$ represents the agent’s belief about the current state given the history of observations and actions. Learning these models enables the agent to construct a comprehensive world model that can be used for planning and decision-making in complex, partially observable environments.
Incorporating world models into MARL based on generative EmCom is natural and reasonable. By estimating $z_{t}$ using the inference model of a world model $q(z_{t}|x_{1:t},a_{1:t-1})$, an agent can estimate its state. Agents are expected to perform EmCom in the same manner as described in Section 3.1. Figure 4 shows a PGM of generative EmCom on POMDPs. This demonstrates the theoretical connection between EmCom or language and world models. When considered in this way, CPC is the theory explaining that language is something that integrates world models of populated agents. This will be discussed in more detail in Section 5.
Figure 4: Graphical model of generative EmCom involving world models.
Several studies linking EmCom and world models have been conducted. Recent studies on EmLang have focused on MARL approaches, aiming to autonomously develop languages to solve tasks [2]. Because each agent typically has access only to its own observations, inferring the underlying state of the environment becomes crucial. Consequently, researchers have explored enabling RL agents to develop EmLangs for inter-agent communication while simultaneously learning world models to infer environmental states. Notably, not all MARL communication studies focus on developing symbol-based languages; some treat communication as simple information exchange. This distinction separates “ EmLang” from “learning tasks with communication.” [49]
Alexander et al. proposed a message-conditioned world model that improves the quality of communication among agents, facilitating the emergence of meaningful messages and enhancing the interpretability of the model [84]. Similarly, Kenzo Lobos-Tsunekawa et al. extended the widely used Dreamer world model by introducing a global world model, used exclusively during training. By incorporating information from all agents, they achieved improved sample efficiency through shared imagination [85]. Toledo et al. also used Dreamer but introduced GNNs to facilitate communication; however, their work is classified under ”learning tasks with communication” [86] because it does not aim to develop a model of EmLang.
IV Relationship with Conventional EmCom
Recent years have witnessed an increasing presence of research papers focused on EmCom at ML conferences such as NeurIPS and ICLR. This trend underscores the renewed interest in EmCom. In the field of EmCom, Lewis’ signaling game [87] framework is often used as a model of communication while other various formulations have also been proposed [35, 36, 37]. The game players, or agents, are optimized to achieve some goals, resulting in the communication protocols, which are referred to as EmLang. In the EmCom field, the extent to which EmLangs resemble human languages is frequently questioned from the perspectives of evolutionary and/or computational linguistics. For instance, questions have been raised about whether EmLangs have compositionality [88, 89] or whether they follow the well-known statistical universal properties of natural languages [90, 91, 92, 93]. Additionally, cases exist where cognitive constraints were modeled to induce characteristics of natural languages [94, 95].
The signaling game is a simple communication model that involves only a sender $S_{\boldsymbol{\phi}}(m|x)$ and a receiver $R_{\boldsymbol{\theta}}(x|m)$ and only allows unidirectional communication from the sender to the receiver. At each play, the signaling game proceeds as follows:
  1. 1.
    Observation: Sender $S_{\boldsymbol{\phi}}$ obtains an observation $x$, that is, $x\sim P(x)$.
  2. 2.
    Signaling: Sender $S_{\boldsymbol{\phi}}$ generates a message $m$ from the observation $x$, that is, $m\sim S_{\boldsymbol{\phi}}(m|x)$.
  3. 3.
    Reconstruction: Receiver $R_{\boldsymbol{\theta}}$ attempts to reconstruct the original observation $x$ from the message $m$ via $R_{\boldsymbol{\theta}}(x|m)$.
Sender $S_{\boldsymbol{\phi}}$ and receiver $R_{\boldsymbol{\theta}}$ are optimized via a gradient-based method toward successful communication. Conventionally, the objective function of the signaling game (to be maximized) is defined as follows [96, 97]:
$\mathcal{J}_{\textup{MI}}(\boldsymbol{\phi},\boldsymbol{\theta})\mathrel{:=}\mathbb{E}_{P(x),S_{\boldsymbol{\phi}}(m|x)}[\log R_{\boldsymbol{\theta}}(x|m)].$ (12)
We refer to $\mathcal{J}_{\textup{MI}}$ as the MI-maximizing objective function because it is known to be a variational lower bound (up to constant) of the following mutual information between $X$ and $M$ [98, 99]:
$I_{\boldsymbol{\phi}}(X;M)\mathrel{:=}\mathbb{E}_{P(x),S_{\boldsymbol{\phi}}(m|x)}\left[\log\frac{S_{\boldsymbol{\phi}}(m|x)}{\mathbb{E}_{P(x^{\prime})}[S_{\boldsymbol{\phi}}(m|x^{\prime})]}\right].$ (13)
This implies that the conventional signaling game in the field of EmCom has been formulated as a problem of maximizing the mutual information between $X$ and $M$, where $X$ and $M$ denote random variables corresponding to the realizations $x$ and $m$, respectively. In contrast, a recent study [100] has proposed a reformulation of the signaling game as a problem of maximizing the evidence lower bound (ELBO):
[…]
We refer to $\mathcal{J}_{\textup{ELBO}}$ as the ELBO-maximizing objective function, contrasting it with the MI-maximizing objective $\mathcal{J}_{\textup{MI}}$. By adopting the ELBO-maximizing objective function, we can introduce concepts from computational psycholinguistics into signaling games. To observe this, let us transform the ELBO maximizing objective function as follows:
$\begin{split}\mathcal{J}_{\textup{ELBO}}(\boldsymbol{\phi},\boldsymbol{\theta})&\mathrel{:=}\mathbb{E}_{P(x),S_{\boldsymbol{\phi}}(m|x)}[\underbrace{\log R_{\theta}(x|m)}_{\textup{communication}}+\beta\underbrace{\log P_{\theta}(m)}_{\begin{subarray}{c}\textup{(negative)}\\ \textup{surprisal}\end{subarray}}]\\ &\phantom{\mathrel{:=}}-\beta\underbrace{\mathbb{E}_{P(x)}\mathcal{H}(S_{\phi}(M|x))}_{\begin{subarray}{c}\textup{entropy}\\ \textup{maximizer}\end{subarray}},\end{split}$ (15)
Here, a term known as surprisal appears, which is a concept commonly used in computational psycholinguistics [101, 102, 103]. Surprisal is assumed to represent the cognitive load experienced by a listener/reader (or the receiver in the signaling game) when processing a sentence. Therefore, the ELBO-maximizing signaling game naturally models the trade-off between information transmission and surprisal.
A similar discussion involves modeling the trade-off between information transmission and efficiency, where studies have used the (variational) information bottleneck (IB, VIB) framework to model communication [104, 105, 106]. In fact, VIB is proven to be a generalization of (beta-)VAE [107, 108], resulting in similar models. In addition, a contemporary work [109] shows that a variant framework known as a referential game can also be reformulated with an ELBO-like objective, analogous in structure to that of a conditional VAE [110, 111].
In relation to discussions in the field of evolutionary linguistics, some studies have also incorporated ILM into the EmCom framework [112], which is another important research theme. The ILM is a framework that models generational changes, where supervised learning is repeatedly performed from parent agents to child agents. In the context of VAE and VIB, however, little discussion exists on modeling generational changes, and this remains a future challenge when considering generative symbol emergence.
The formulation of the signaling game presented in this section, that is, generative model-based re-formulation of conventional EmCom, shares some fundamental connections with the generative EmCom and CPC frameworks discussed in other sections. All these formulations can be interpreted as representation learning with messages (signals, signs, or sentences) serving as latent variables in generative models when using ELBO-type (or VIB-type) objectives. However, notable differences exist: Although the CPC hypothesis does not specify a particular method for posterior distribution estimation, the signaling game explicitly adopts amortized variational inference. Furthermore, the signaling game assumes asymmetric roles between senders and receivers, which is not necessarily the case in the CPC frameworks. Additionally, the extension of ELBO-type formulation to populated settings, where multiple agents interact, remains an open challenge in the signaling game framework and requires further investigation [^1] An ELBO-type formulation based on the FEP for populated settings of generative EmCom is described in [45]. However, an inference procedure, that is, a language game, for enabling agents to let a symbol system emerge has not been described..
V Large Language Model as Collective World models
Figure 5: Three levels of the relationship between language, perceptual, and action information. Left: (A) Image captioning and generation tasks corresponding language to a still image. (B) Video captioning and generation tasks corresponding language to a video, that is, a sequence of visual stimuli. (C) Action-dependent video captioning and generation corresponding language to dynamic perceptual and action information, which corresponds to a world model and a policy conditioned by language. Right: probabilistic generative models corresponding to each of (A) – (C).
V-A Our argument
Based on the discussion in Sections 2 and 3, we explain the main argument of this paper, stating that LLMs are collective world models.
To clarify the link between this argument and existing studies relating language and perceptual information, Figure 5 shows three levels of tasks relating language to visual information (an example of multimodal sensory information) and action streams, along with PGMs for each. These represent (A) image captioning and generation tasks, corresponding language to a still image [113, 114, 115]; (B) video captioning and generation tasks, corresponding language to a video, that is, a sequence of visual stimuli [116, 117]; and (C) action-dependent video captioning and generation, corresponding language to dynamic perceptual and action information, respectively [^2] Note that we redefine the generative process of world models assuming that the policies of agents that generate $a_{t}$ from $z_{t}$ are also a part of the generative process of this world model.. In particular, from the perspective of representation learning, (A) and (B) correspond to representation learning of images and videos, whereas (C) corresponds to world models explained in Section III-B. Language-conditioned world models and robotics foundation models, known as vision-language-action models, studied in robotics and autonomous vehicles correspond to (C) [118, 119, 120, 121, 122].
Corresponding to the three levels of captioning and generation tasks, we can consider generative EmCom. Figure 5 describes PGMs for generative EmCom corresponding to the three levels, respectively. Notably, they are only an instance of the generative model shown in the CPC hypothesis (Figure 1) and generalization of studies based on the MHNG described in Sections 2 and 3.
Figure 6: PGMs for generative EmCom corresponding to three levels of complexity: (A) image-, (B) video-, and (C) action- and video-based tasks. These models represent instances of the CPC hypothesis and generalizations of MHNG studies.
Here, $\theta^{k}$ represents the global parameters of the $k$-th agent, for instance, the parameters of neural networks for representation learning. In (C), the global parameter $\bar{\theta}^{k}$ includes not only the parameters of the prediction model of the world (i.e., Equations 9 and 10 ) but also those of policies $p(a_{t}|z_{t})$. Let us consider (C) as a generative model corresponding to a single super-agent, virtually. The super-agent has multimodal sensory information combining the observations $\{x^{k}_{1:t}\}_{k}$ and actions $\{a^{k}_{1:t}\}_{k}$ of every agent. In this case, (C) can be considered as a type of hierarchical world model combining $K$ agents and integrating their experiences. We can refer to the world model owned by the super-agent as a collective world model, where $\{\phi,\{\theta^{k}\}_{k}\}$ are its parameters.
From this perspective, the corpus data that comprise sentences uttered by agents are considered as a sample from approximate posterior distributions $m_{j}^{[i]}\sim q(m|\{x^{k}_{j,1:t},a^{k}_{j,1:t}\}_{k})$, where $m_{j}^{[i]}$ denotes the $i$-th sentence describing the $j$-th observation $\{x^{k}_{j,1:t}\}_{k}$ and action $\{a^{k}_{j,1:t}\}_{k}$.
Language modeling implies approximating the distribution of word sequences $p(m|\phi^{k})$ in a corpus, where $\phi^{k}$ denotes a parameter of the language model of the $k$-th agent. This implies that the language model approximates the posterior distribution $q(m|\{x^{k}_{j,1:t},a^{k}_{j,1:t}\}_{k})\approx p(m|\theta^{LM})$.
The rationale behind distributional semantics as modeled by LLMs can be comprehended through the CPC framework. When language emerges through CPC, it functions as a latent space that develops through representation learning based on the collective experiences of all agents — in this case, humans. These experiences encompass not only external sensory information but also interception and behavioral decision-making. Within this framework, sentences can be mathematically considered as samples drawn from the posterior distribution over messages $m$, formally expressed as $q(m|\{x^{k},a^{k}\}_{k})$, where $x^{k}$ and $a^{k}$ represent the observations and actions of agent $k$, respectively. This formulation naturally leads to the structural reflection of the distribution of $\{x^{k},a^{k}\}_{k}$ in that of $m$, analogous to the working mechanism of representation learning in individual agents.
A key insight is that language serves as a shared external representation that effectively combines the world models of multiple human agents. When examining the graphical model depicted in Figure 6 (C) as a cognitive model of a single ”super-agent,” we can interpret it as a factorized hierarchical world model incorporating $K$ distinct modalities. Through this lens, an LLM that models the distribution $q(m|\{x^{k},a^{k}\}_{k})$ can be considered as a collective world model — one that integrates the diverse experiences and knowledge of multiple agents into a unified representational space. This provides a theoretical foundation for understanding why LLMs trained on large-scale human-generated text corpora appear to capture rich world knowledge.
The emergence of these phenomena can be comprehensively understood through the lens of the CPC hypothesis. The hypothesis suggests that language evolution occurs through a process of CPC, where the symbol system (language) emerges to maximize the predictability of multimodal sensory-motor information obtained by members of a society. This theoretical framework helps explain why LLMs, trained on the linguistic outputs of this process, inherit structural properties that reflect collective human knowledge and experience.
The CPC hypothesis thus provides a principled explanation for the acquisition of world knowledge by language models without direct sensorimotor experience by tapping into the accumulated wisdom encoded in human language through this collective process.
V-B Existing argument
LLMs learn to predict text tokens in an autoregressive manner, and unlike human cognition or prediction, they are not designed to interact with environments or achieve specific goals beyond word prediction. Nevertheless, some studies have suggested that LLMs may implicitly contain a world model that simulates our physical world.
Hao et al. proposed “reasoning via planning,” where an LLM is reused as a world model to perform logical reasoning, employing Monte Carlo tree search to reason in a more planned manner [123]. This method demonstrated higher performance compared to the existing chain-of-thought approach. Li et al. used a GPT model trained only on Othello game logs and confirmed that the model made predictions based on an internal world model by performing tests using explicit intervention operations that changed the activation of the middle layer of the model [124]. This suggests that LLM does not simply rely on statistical inference but constructs some degree of world model in systems such as games. In contrast, Wang et al. developed a benchmark involving text games to test whether transition models or policies can be derived from language models and revealed that LLMs are not reliable world simulators [125].
Although these studies investigate whether LLMs can be used as world models, studies examining whether LLMs inherently possess such models are underway. Gurnee et al. demonstrated that LLMs acquire spatiotemporal linear representations of the real world, specifically identifying individual spatial and temporal neurons [21]. The authors argue that this corresponds to the basic elements of a world model, although they did not show the acquisition of a causal model that includes actions. Andreas argues that during the process of predicting the next token, a language model infers the internal states, such as the beliefs, desires, and intentions of the speaker, from the context and reflects them in text generation or speech directed at other agents, suggesting the presence of a world model within the linguistic space [126]. Previous studies also indicated the possibility of representing belief states and intentions during text modeling. However, this pertains to agents within the linguistic space and does not explain connections with data derived from the physical world.
[…]
VI Conclusion
This study proposed a theoretical framework that unified EmCom, world models, and LLMs through the lens of CPC. We introduced the concept of generative EmCom as an alternative formulation of the conventional EmCom, which is based on a discriminative model-based language game such as referential games, and described their relationships. The concept, generative EmCom, is based on the CPC hypothesis, which demonstrates the modeling of language emergence as decentralized Bayesian inference of shared latent representations. We showed the application of this framework to MARL and provided a novel perspective on LLMs as collective world models that integrate the diverse experiences and knowledge of multiple agents into a unified representational space.
The theoretical framework proposed here offers several important insights. First, it provides a principled explanation for how language models can acquire world knowledge without direct sensorimotor experience, by tapping into the accumulated wisdom encoded in human language through CPC. Second, it bridges the gap between individual cognitive development and collective language evolution by showing how both can be comprehended through the lens of representation learning and free energy minimization. Third, it suggests that the success of LLMs may be fundamentally linked to their ability to capture and integrate the collective world models acquired by humans, i.e., through embodied sensorimotor interactions with the world.
However, several limitations of the current work should be addressed. First, although we provide a theoretical framework, direct empirical evidence for the collective world model hypothesis remains limited. The relationship between neural representations in language models and human conceptual structures needs further investigation. Second, the proposed framework primarily focuses on the emergence of linguistic structure and meaning but does not fully address the emergence of pragmatic aspects of language use. Third, the current formulation may not fully capture the dynamic, interactive nature of human language evolution, for instance, language evolution over generations and through the interaction of several agents in an open world.
From the perspective of world models, understanding the influence of collective world models on the environmental adaptation of agents is also important. According to the proposed theoretical framework, a language formed by multiple agents with the same embodiment learning in the same environment should serve as an appropriate prior distribution for individual world models. In other words, EmLang should accelerate both world model learning and following environmental adaptation based on RL and other methods. Obtaining constructive evidence for this relationship is also crucial.
Our argument and theoretical framework initiate several promising directions for future studies. Although we have established the basic connections between emergent communication, world models, and LLMs, significant work remains to validate and extend these ideas. Key priorities include developing experimental paradigms to test the CPC hypothesis, creating more sophisticated computational models and implementations of generative EmCom, and investigating how language formed by embodied agents sharing similar environments could serve as an effective prior distribution for world model learning. From a theoretical perspective, a deeper mathematical formalization of the interaction between individual and collective learning processes will be crucial. In addition, this framework has important practical implications for the development of more capable multi-agent systems, improved human–AI interaction, and embodied AIs that rapidly adapt to their physical worlds and communicate their knowledge using EmLang. By pursuing these directions, we can work toward a more comprehensive understanding of language emergence and its role in environmental adaptation while advancing both the theoretical foundations and practical applications of AI systems.